Could you help me please
f(x)=4x-5 and g(x)=-3x+2, find f(g(-2))
thank you for your time

Answer by kapilsinghi(68)

f(x)=4x-5 and g(x)=-3x+2, find f(g(-2))
g(-2) = -3(-2) + 2
= 6+2
= 8
f(g(-2)) = f(8)
= 4(8) - 5
= 32 - 5
= 27
Regards,
Kapil
You can send your questions directely at kapilsinghi123@gmail.com